OpenStudy (anonymous):

Which of the following expressions is a factor of y^7 - 128 A) y - 1 B) y-3 C) y-2 D) y+3

OpenStudy (campbell_st):

well 2^7 = 128 so perhaps y - 2 becuase P(y) = y^7 - 128 P(2) = 2^7 - 128 = 0 using the factor theorem